At a summer camp, each girl is assigned to one of 48 cabins that in total can accommodate 344 campers. There are two types of cabins-small ones that can house 6 girls and large ones that can house 10 girls. Last summer even though 344 girls had registered for the camp, only 342 could come. So one of the large cabins had only 8 girls. How many more small cabins are there than large cabins?
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

6S+10L=344
S+L=48;L=48-S
6S+480-10S=344
-4S=-136
S=34 cabins with 6 (204 capacity)
L=14 cabins with 10 (140 capacity)
The answer is 20 more small cabins

Answer by ikleyn(52777)

.
At a summer camp, each girl is assigned to one of 48 cabins that in total can accommodate 344 campers.
There are two types of cabins-small ones that can house 6 girls and large ones that can house 10 girls.
Last summer even though 344 girls had registered for the camp, only 342 could come.
So one of the large cabins had only 8 girls. How many more small cabins are there than large cabins?
~~~~~~~~~~~~~~~

L = the number of large cabins.

(48-L) = the number of small cabins.


From the condition, you have this equation


    10*(L-1) + 8 + 6*(48-L) = 342


Simplify and find L


    10L - 10 + 8 + 6*48 - 6L = 342

    4L = 342 + 10 - 8 - 6*48

    4L = 56

     L = 56/4 = 14.


There were 14 large cabins and 48-14 = 34 small cabins.

The number of small cabins was 20 more that the number of large cabins.    ANSWER


CHECK.  10*(14-1) + 8 + 6*34 = 342 girls, in total.    ! Correct !

Solved.